OpenStudy (anonymous):

find a simplified form of the expression. give your answer in scientific notation (3 x 10^8)x(9x10^7)

OpenStudy (shinalcantara):

you just have to multiply the bases. 3 x 9 = ? add the exponents of the base 10 10^8(10^7)= 10^(8+7)

OpenStudy (anonymous):

thats it ?

OpenStudy (shinalcantara):

You will have it as \[27 \times 10^{15}\] But since you are to have it in scientific notation and the rule to expressing numbers in scientific notation states that the factor to be multiplied by 10^n is more than or equal to 1 but less than 10. Then you'll have it as \[2.7 \times 10^{16}\]

OpenStudy (anonymous):

can you help me with another one

OpenStudy (perl):